The Alpine House, Inc., is a large retailer of snow skis. The company assembled the information shown below for the quarter ended March 31: Amount Sales Selling price per pair of Variable selling expense per pair of skis Variable administrative expense per pair of skis Total fixed selling expense Total fixed administrative expense Beginning merchandise inventory Ending merchandise inventory Merchandise purchases s1,276,000 440 47 20 $130,000 $ 120,000 65,000 105,000 $ 290,000 skis Required: 1. Prepare a traditional income statement for the quarter ended March 31. 2. Prepare a contribution format income statement for the quarter ended March 31 3. What was the contribution margin per unit? Complete this question by entering your answers in the tabs below Required 1Required 2 Required 3 Prepare a traditional income statement for the quarter ended March 31 The Alpine House, Inc. Traditional Income Statement Selling and administrative expenses:Complete this question by entering your answers in the tabs below Required 1Required 2 Required 3 Prepare a traditional income statement for the quarter ended March 31 The Alpine House, Inc Traditional Income Statement Selling and administrative expenses 03 Required 1 Required 2>Required 1Required 2Required 3 Prepare a contribution format income statement for the quarter ended The Alpine House, Indc Contribution Format Income Statement Variable expenses: 0 Fixed expenses: 0March 31: Amount Sales Selling price per pair of skis Variable selling expense per pair of skis Variable administrative expense per pair of skis Total fixed selling expense Total fixed administrative expense Beginning merchandise inventory Ending merchandise inventory Merchandise purchases $1,276,000 440 47 20 $ 130,000 120,000 $65, 000 105,000 290,000 Required: 1. Prepare a traditional income statement for the quarter ended March 31. 2. Prepare a contribution format income statement for the quarter ended March 31 3. What was the contribution margin per unit? Complete this question by entering your answers in the tabs below Required 1Required 2 Required 3 What was the contribution margin per unit? (Round your final answer to nearest whole dollar.) Contribution margin per unit K Required 2 Required 3
Answer #1

Please find below the answer -

Answer 1_
Unit sold = 1276000/440 2900
The Alpine House Inc.
Traditional Income Statement
Sales 1276000
Less: Cost of Goods Sold 250000
Gross Margin 1026000
Less: Selling and Administrative Expenses:
Selling Expenses 266300
Administrative Expenses 178000 444300
Net Operating Income 581700
Working
Cost of goods sold = Beginning inventory + Purchase - Closing inventory 250000
=65000+290000-105000
Selling Expenses = =130000+2900*47 266300
Administrative Expenses = =120000+2900*20 178000
Answer 2_
The Alpine House Inc.
Contribution Format Income Statement
Sales 1276000
Less: Variable Expenses:
Cost of Goods Sold 250000
Selling Expenses 136300
Administrative Expenses 58000 444300
Contribution Margin 831700
Less: Fixed Expenses:
Selling Expenses 130000
Administrative Expenses 120000 250000
Net Operating Income 581700
Answer 3_
Contribution margin per unit = Total contribution / Number of unit =831700/2900
                        287